OpenStudy (anonymous):

the integral of 4/(x^2 +4) from -10 to 10

OpenStudy (anonymous):

\[4 \int\limits_{-10}^{10} \frac{dx}{x^2+4}=\frac{4}{2}\arctan(\frac{x}{2})|_{-10}^{10}\] \[2(\arctan(5)-\arctan(-5)) \rightarrow 2(\arctan(5)+\arctan(5))=4\arctan(5)\]

OpenStudy (anonymous):

hi!

OpenStudy (anonymous):

I reposted this to see if anyone had a different answer than the one we had yesterday

OpenStudy (anonymous):

b/c i think that is the right answer, but the program says it is wrong

OpenStudy (anonymous):

I don't know what to tell you. Other than e-mail your instructor. :/

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

The graph is not bounded by -10/10, it's actually infinity both ways, because the graph is unbounded on the x-axis.

OpenStudy (anonymous):

It doesn't matter though.